OpenStudy (anonymous):

First one to answer and explains gets a medal!solve the equation? I 3x + 2 I =5

OpenStudy (anonymous):

just to confirm you have two variables? I and x?

OpenStudy (anonymous):

no its not an I is a abs

OpenStudy (anonymous):

Oh ok then you know that IxI you know that x could have been -x or +x. so we will apply the same concept here: either 3x+2 = 5 OR -(3x+2)=5 then you will have two values for X :)

OpenStudy (anonymous):

what so what is the answer

OpenStudy (anonymous):

(3x+2) = 5 --> x =1 (3x+2) =-5 --> x= -7/3

OpenStudy (anonymous):

and how did you get that?

OpenStudy (anonymous):

magic ;)

OpenStudy (anonymous):

no forreal

OpenStudy (anonymous):

Oh ok then you know that IxI you know that x could have been -x or +x. so we will apply the same concept here: either 3x+2 = 5 OR -(3x+2)=5 then you will have two values for X :) 3x=5-2 -3x-2=5 3x=3 -3x=5+2 x=1 -3x=7 >>>>> x= -7/3